# LocoVR: Multiuser Indoor Locomotion Dataset in Virtual Reality

### v1

@@ Abstract
Understanding human locomotion is crucial for AI agents such as robots, particularly in complex indoor home environments. Modeling human trajectories in these spaces requires insight into how individuals maneuver around physical obstacles and manage social navigation dynamics. These dynamics include subtle behaviors influenced by proxemics - the social use of space, such as stepping aside to allow others to pass or choosing longer routes to avoid collisions. Previous research has developed datasets of human motion in indoor scenes, but these are often limited in scale and lack the nuanced social navigation dynamics common in home environments. To address this, we present LocoVR, a dataset of 7000+ two-person trajectories captured in virtual reality from over 130 different indoor home environments. LocoVR provides full body pose data and precise spatial information, along with rich examples of socially-motivated movement behaviors. For example, the dataset captures instances of individuals navigating around each other in narrow spaces, adjusting paths to respect personal boundaries in living areas, and coordinating movements in high-traffic zones like entryways and kitchens. Our evaluation shows that LocoVR significantly enhances model performance in three practical indoor tasks utilizing human trajectories, and demonstrates predicting socially-aware navigation patterns in home environments.
Project page:
https://sites.google.com/view/locovr?usp=sharing
Github:
https://github.com/kojirotakeyama/LocoVR
@@ 1 Introduction
To overcome data collection challenges, we propose LocoVR, a dataset captured in virtual reality (VR) that efficiently captures detailed spatial information, human-scene interactions, and human-human social motion behaviors across diverse indoor environments. LocoVR captures task-focused movements of two people in over 130 home settings, including their full body motions, head orientations, and precise spatial data. Crucially, LocoVR captures motion proxemics - the social use of space, such as yielding in narrow spaces, maintaining personal distances in shared areas, and coordinating movements in high-traffic considering the Interpersonal Adaptation Theory Burgoon et al. (1995). These proxemics-based motion behaviors, often missing in current datasets, serve as a form of non-verbal communication, and are influenced by factors such as the relationship between individuals and their cultural backgrounds Hall (1963); Watson (2014). Human social dynamics can provide valuable insights for home robots to navigate domestic spaces more naturally while adhering to implicit social norms. Prior work has explored proxemics in the context of human-robot interaction Mead & Matarić (2017), proposing a computational framework that considers how a human and a robot experience co-present interaction Mead & Mataric (2012).
@@
We demonstrate the utility of our dataset through three practical applications: global path prediction, trajectory prediction, and goal area prediction (Figure1). The first two tasks showcase the dataset’s capability to facilitate geometrically and socially aware path predictions, while the last task demonstrates its versatility in supporting a broad spectrum of applications. Our key contributions are outlined as follows.
@@
Figure 1: Overview of the data collection VR system, snippets of the LocoVR dataset, and the dataset applications. We collected multi-person trajectories in 131 complex indoor scenes, including full body motions with head orientation, precise geometry, and fully labeled scene semantics. In the dataset applications, predicted and past trajectories are shown in red and blue, respectively.
@@ 2 Related work
2.1 Human trajectory datasets
Existing human trajectory datasets can be categorized as human trajectories in outdoor and indoor scenes. A significant amount of prior work Zhou et al. (2012); Robicquet et al. (2016); Ess et al. (2007); Kothari et al. (2021) has focused on building human trajectory datasets in outdoor scenes. These datasets used aerial cameras to capture human trajectories and surrounding scene images. They have trained machine learning models to learn how humans move by considering other humans’ motions and environmental constraints. However, these approaches are not directly applicable to indoor scenes due to the significant complexity of indoor scenes regarding geometry and semantics and the much smaller scale of indoor environments compared with outdoor environments. For indoor scenes, previous research Rudenko et al. (2020); Schreiter et al. (2024); Vendrow et al. (2023) has collected human trajectories using Lidars. However, the scenes were set in an open public space, similar to outdoor scenes, with incomplete scene data due to the Lidar blind spots. Additionally, these indoor datasets are limited in scale as acquiring scene variation is challenging and time-consuming. LocoVR aims to solve these limitations using a VR system, which enables efficient, rapid, and accurate data collection across diverse indoor scene variations.
2.2 Human motion datasets
@@
Understanding human motion is important for various research problems, such as synthesizing motion for 3D environments. Prior work has studied the 3D human motion problem and built datasets to support research in this domain. One line of research has explored human-scene interactions and focused on understanding how environmental constraints affect human behavior, such as GIMO (Zheng et al., 2022; Guzov et al., 2021; Zhang et al., 2022). A challenge faced by human-scene interaction datasets is capturing scene variation is difficult and expensive. To address this challenge, prior research Wang et al. (2022); Araújo et al. (2023) has worked on synthesizing human motions in virtual environments and game engines or generating human motions with generative AI models. However, these synthesized motions do not necessarily follow human behavior principles, especially when social motion behaviors are involved. Additionally, the existing datasets focus mainly on single-person motions; few datasets have been collected for multi-person interactions. However, understanding multi-person social motion behavior is critical for many real-world use cases, such as robot task assistance and human-robot collaboration. Moreover, existing human motion datasets provide limited locomotion data needed to understand room-scale human motion dynamics in indoor environments, as they primarily focus on capturing fine-grained details of individual actions, such as sitting on a chair or opening a fridge. To fill these research gaps, we introduce LocoVR, a novel dataset comprising two-person trajectories with social motion behaviors, collected using VR to allow a large number of trajectories across diverse indoor scene variations.
@@ 3.1 Overview
Table 1 summarizes the statistics of existing human trajectory and motion datasets and LocoVR. Our dataset contains 2500K frames of human trajectories in 131 scenes (see Figure 2 for examples). The number of scenes surpasses all the real human motion datasets. We collected two-person trajectories that are geometrically and socially aware, which is not included in most of the compared datasets. The number of trajectories is 7071 in total (see Appendix I for detailed statistics of LocoVR). LocoVR facilitates the enhancement of task performances from both geometric and social perspectives in unseen, complex, and confined indoor environments. Also, it includes full-body human poses, along with head orientation data in addition to trajectories. These additional observations can facilitate a deeper understanding of human locomotion and enhance model performance.
@@
$Count$ $Semantics$ $Geometry$ $Location$ $Pos/Pose$ $Multi$ $Motion^{*}$ $Target-action$
HPS
(Guzov et al., 2021)
@@
Grand Station
(Zhou et al., 2012)
@@
THOR-MAGNI
(Schreiter et al., 2024)
@@
2500K 131
✓(3D mesh) Indoor 3D, Head Real Trajectory, Social motion
@@ 4 Evaluation
    Global path prediction: This task estimates a static global path from a starting point to a goal location, which can be used to predict human global paths or plan human-like global paths for robots. Our dataset demonstrates the ability to learn such human-like paths that consider obstacle avoidance, efficiency, and social motion behaviors, such as maintaining social distance when passing or choosing longer routes to avoid collisions. The input includes the past trajectories of two people, $p_{1}$ and $p_{2}$ (length=1.0s, interval=0.067s), past heading directions of $p_{1}$ and $p_{2}$ (length=1.0s, interval=0.067s), scene map, and goal position. The output is a static path from the start to the goal.
@@ 5 Limitations and future work
In this paper, we evaluated three tasks incorporating 2D trajectory and head pose data to demonstrate the effectiveness of our dataset. Future work could utilize LocoVR’s comprehensive spatial data, including 3D geometry, scene semantics, and full-body 3D human poses, for a wider range of applications, such as full-body motion prediction.
Moreover, gaps existed between virtual environments and the real world, such as the lack of physicality in VR objects and lower avatar expressivity compared to real people (See Appendix J for details). Although we have evaluated the prediction models through the data collected in real space (LocoReal), the gaps could affect detailed human behaviors, such as walking speed or interpretation of non-verbal communication through facial expressions. Investigating these impacts and comparing LocoVR to fully synthetic datasets are important research questions for future work.
@@ 6 Conclusion
To model geometrically and socially aware human trajectories in complex indoor environments, we introduced the LocoVR dataset, which captures two-person social motion behaviors across 131 home environments, including full-body motions and detailed spatial information. In the experiments, we introduced three indoor tasks that utilize human trajectory: global path prediction, trajectory prediction, and goal prediction. Experimental results showed that the models trained with LocoVR outperformed other prior indoor datasets evaluated on the real-world test data. This indicates that our dataset facilitates adaptation to unseen indoor environments with complex geometries and social motion behaviors across a variety of tasks. Furthermore, these findings demonstrate the potential of virtual environments for training models that generalize well to real-world applications. We envision the data collection method to expand the variety of indoor scenes used for training and propose the experiments as a standard benchmark for future research on human motion and trajectory analysis in indoor settings.
Reproducibility statement
We performed three trials on the training and testing process to ensure the reproducibility of the evaluations shown in the paper. Table 2,3,4 report averaged ADEs over three trials $\pm$ standard deviations.
@@ Appendix C Ablation study
Appendix D Testing on GIMO
While the main paper utilized real-world trajectory data (LocoReal) as the test data, we conducted an additional experiment to further validate the contribution of LocoVR using an alternative test dataset. Given that THOR-MAGNI has fewer variational scenes, we selected GIMO as the test data. In this experiment, GIMO was divided into 70% for training, 15% for validation, and 15% for testing, ensuring that the scenes in the training, validation, and test sets were mutually exclusive. Additionally, to mitigate potential bias in the test data, we performed five random splits to generate five different datasets and averaged the results. Figures 8 through 10 illustrate the evaluation results for global path planning, trajectory prediction, and goal prediction, respectively. The evaluation results demonstrate that the model trained with LocoVR consistently outperforms those trained on other datasets. This superior performance is attributed to the enhanced generalization capabilities provided by LocoVR’s extensive coverage of scenes and trajectories.
@@ Appendix H Data augmentation
For the global path prediction, the time window with a length of 7.0s (1.0s for past trajectory and 6.0s for future trajectory) slides with an interval of 0.67s to extract trajectory segments from the raw trajectory. A part of the trajectory segment that is out of the raw trajectory is padded with the last value observed in the raw trajectory.
For the trajectory prediction, the time window with a length of 4.0s (1.0s for past trajectory and 3.0s for future trajectory) slides with an interval of 0.67s to extract trajectory segments from the raw trajectory. All the trajectory segments are within the raw trajectory because the task focuses on trajectory prediction in local areas on the way to the goal.
For the goal prediction, the time window with a length of 6.0s (for past trajectory) slides with an interval of 0.67s to extract trajectory segments from the raw trajectory. A part of the trajectory segment that is out of the raw trajectory is padded with the initial value observed in the raw trajectory.

### v2

@@ Abstract
Understanding human locomotion is crucial for AI agents such as robots, particularly in complex indoor home environments. Modeling human trajectories in these spaces requires insight into how individuals maneuver around physical obstacles and manage social navigation dynamics. These dynamics include subtle behaviors influenced by proxemics - the social use of space, such as stepping aside to allow others to pass or choosing longer routes to avoid collisions. Previous research has developed datasets of human motion in indoor scenes, but these are often limited in scale and lack the nuanced social navigation dynamics common in home environments. To address this, we present LocoVR, a dataset of 7000+ two-person trajectories captured in virtual reality from over 130 different indoor home environments. LocoVR provides accurate trajectory data and precise spatial information, along with rich examples of socially-motivated movement behaviors. For example, the dataset captures instances of individuals navigating around each other in narrow spaces, adjusting paths to respect personal boundaries in living areas, and coordinating movements in high-traffic zones like entryways and kitchens. Our evaluation shows that LocoVR significantly enhances model performance in three practical indoor tasks utilizing human trajectories, and demonstrates predicting socially-aware navigation patterns in home environments. The dataset and evaluation code are available at https://github.com/kt2024-hal/LocoVR.
@@ 1 Introduction
To overcome data collection challenges, we propose LocoVR, a dataset captured in virtual reality (VR) that efficiently captures detailed spatial information, human-scene interactions, and human-human social motion behaviors across diverse indoor environments. LocoVR captures task-focused movements of two people in over 130 home settings, including their trajectories, head orientations, and precise spatial data. Crucially, LocoVR captures motion proxemics - the social use of space, such as yielding in narrow spaces, maintaining personal distances in shared areas, and coordinating movements in high-traffic considering the Interpersonal Adaptation Theory Burgoon et al. (1995). These proxemics-based motion behaviors, often missing in current datasets, serve as a form of non-verbal communication, and are influenced by factors such as the relationship between individuals and their cultural backgrounds Hall (1963); Watson (2014). Human social dynamics can provide valuable insights for home robots to navigate domestic spaces more naturally while adhering to implicit social norms.
@@
We demonstrate our dataset through three tasks: global path prediction, trajectory prediction, and goal area prediction (Figure1). The first two tasks showcase the dataset’s capability to facilitate geometrically and socially aware path predictions, while the last task demonstrates its versatility in supporting a broad spectrum of applications. Our key contributions are outlined as follows.
@@
Figure 1: Overview of the data collection VR system, snippets of the LocoVR dataset, and the dataset applications. We collected multi-person trajectories in 131 complex indoor scenes, including trajectory with head orientation, precise geometry. In the dataset applications, predicted and past trajectories are shown in red and blue, respectively.
@@ 2 Related work
2.1 Datasets for social navigation
Datasets for social robot navigation can be categorized based on their scene types and the nature of the robot or human agent involved. Studies such as Rudenko et al. (2020); Schreiter et al. (2024); Finean et al. (2023); Manso et al. (2020) provide datasets collected in controlled lab settings where robots coexist with humans. While these datasets offer valuable insights into human-robot interactions in indoor settings, their scene structure is simplistic, and the number of unique scenes is limited, which reduces their capability on diverse real-world indoor settings. On the other hand, datasets like SIT Bae et al. (2024), SACSON Hirose et al. (2023), JRDB Vendrow et al. (2023), and Socnav1 Manso et al. (2020) capture interactions in crowded public spaces by robots or human equipped with cameras and Lidar. Despite their richness in capturing diverse interactions, the open and unstructured nature of these scenes poses challenges for generalizing the findings to indoor settings. Furthermore, the partial observation of scene structures limits their application for indoor navigation tasks, such as global path planning. Datasets such as those proposed by Zhou et al. Zhou et al. (2012), Robicquet et al. Robicquet et al. (2016), Ess et al. Ess et al. (2007), and Kothari et al. Kothari et al. (2021) used aerial cameras to capture human trajectories and surrounding scene images. They provide complementary data on human trajectories and scene geometry across a wide range of environments, enabling the study of crowd dynamics. However, their applicability to indoor scenes is limited due to differences in scene structure, spatial constraints, and interaction dynamics within enclosed spaces.
2.2 Datasets for human motion synthesis and generation
@@
Understanding human motion is important for various research problems, such as synthesizing motion for 3D environments. Prior work has studied the 3D human motion problem and built datasets to support research in this domain. One line of research has explored human-scene interactions and focused on understanding how environmental constraints affect human behavior, such as GIMO (Zheng et al., 2022; Guzov et al., 2021; Zhang et al., 2022). A challenge faced by human-scene interaction datasets is capturing scene variation is difficult and expensive. To address this challenge, prior research Wang et al. (2022); Cao et al. (2020) has worked on synthesizing human motions in virtual environments and game engines or generating human motions with generative AI models. However, these synthesized motions do not necessarily follow human behavior principles, especially when social motion behaviors are involved. Additionally, the existing datasets focus mainly on single-person motions though understanding multi-person social motion behavior is critical for many practical applications, such as human-robot collaboration. Moreover, existing human motion datasets provide limited locomotion data needed to understand room-scale human motion dynamics, as they primarily focus on capturing fine-grained details of individual actions, such as opening a fridge. To fill these research gaps, we introduce LocoVR, a novel dataset comprising two-person trajectories with social motion behaviors, collected using VR to allow a large number of trajectories across diverse indoor scene variations.
2.3 VR systems for human motion analysis
VR technologies including motion tracking systems have been extensively used to study human-scene interactions and behavior across various domainsLee et al. (2023); Takahashi et al. (2021); Simeone et al. (2017); Brookes et al. (2020), including sports strategy evaluation Wang et al. (2024b), road crossing studies Gallo et al. (2024), and navigation in crowded environments Yun et al. (2024). These applications have demonstrated their versatility in replicating real-world scenarios while ensuring safety and repeatability. Datasets specifically developed for VR-based studies focus on distinct scenarios, such as road crossing Wu et al. (2023) and object-approaching tasks Araújo et al. (2023). While these datasets facilitate task-specific behavioral analyses, they lack representation of complex indoor locomotion scenarios involving intricate geometry and multi-person interactions. To address this gap, we introduce a VR system that delivers an immersive locomotion experience across diverse indoor scenes, offering a novel dataset and valuable insights to advance research on human locomotion and interactions in complex indoor environments.
@@ 3.1 Overview
Table 1 summarizes the statistics of existing human trajectory and motion datasets and LocoVR. Our dataset contains 2500K frames of human trajectories in 131 scenes (see Figure 2 for examples). The number of scenes surpasses all the real human motion datasets. We collected two-person trajectories that are geometrically and socially aware, which is not included in most of the compared datasets. The number of trajectories is 7071 in total (see Appendix I for detailed statistics of LocoVR). LocoVR facilitates the enhancement of task performances from both geometric and social perspectives in unseen, complex, and confined indoor environments. Also, it includes body tracker data on head/waist/hands/feet as auxiliary information. These additional observations could facilitate a deeper understanding of human locomotion and enhance model performance.
@@
$Count$ $Geometry$ $Location$ $Pos/Pose$ $Multi$ $Motion^{*}$ $Target-action$
HPS (Guzov et al., 2021)
@@
Grand Station (Zhou et al., 2012)
@@
THOR-MAGNI (Schreiter et al., 2024)
@@
2500K 131 ✓(3D mesh) Indoor 3D, Head Real Trajectory, Social motion
@@ 4 Evaluation
    Global path prediction: This task estimates a static global path from a start to a goal, applicable to goal-conditioned human path prediction or global path planning for robots. The task demonstrates the capability of our dataset to learn physically and socially plausible human trajectory, including social motion behaviors such as maintaining social distance when passing or choosing longer routes to avoid others. The input includes the past trajectories of two people, $p_{1}$ and $p_{2}$ (length=1.0s, interval=0.067s), past heading directions of $p_{1}$ and $p_{2}$ (length=1.0s, interval=0.067s), scene map, and goal position. The output is a static path from the start to the goal.
@@ 5 Limitations and future work
Although we have evaluated the prediction models through the data collected in real space (LocoReal), the gaps could affect detailed human behaviors, such as walking speed or interpretation of non-verbal communication through facial expressions. Investigating these impacts and comparing LocoVR to fully synthetic datasets are important research questions for future work.
While LocoVR is designed to facilitate research on indoor human trajectories, it serves as a foundational resource for exploring the relationships between motion patterns, goal positions, and indoor scene geometries. Beyond its core purpose, LocoVR shows promise for extended tasks such as inferring indoor layouts from trajectory data or predicting human actions by incorporating trajectory Takeyama et al. (2024). Furthermore, it has potential applications in diverse fields, including robot navigation, AI agents in VR, avatar control in both real-worldSchwarz et al. (2023) and VR environmentsWang et al. (2024a), and human action predictionTakeyama et al. (2024). With its versatility, LocoVR is poised to drive innovation across these domains and beyond.
@@ 6 Conclusion
To model geometrically and socially aware human trajectories in complex indoor environments, we introduced the LocoVR dataset, which captures two-person social motion behaviors across 131 home environments, including accurate trajectory and detailed spatial information. In the experiments, we introduced three indoor tasks that utilize human trajectory: global path prediction, trajectory prediction, and goal prediction. Experimental results showed that the models trained with LocoVR outperformed other prior indoor datasets evaluated on the real-world test data. This indicates that our dataset facilitates adaptation to unseen indoor environments with complex geometries and social motion behaviors across a variety of tasks. Furthermore, these findings demonstrate the potential of virtual environments for training models that generalize well to real-world applications. We envision the data collection method to expand the variety of indoor scenes used for training and propose the experiments as a standard benchmark for future research on human motion and trajectory analysis in indoor settings.
Acknowledgment
This research was funded by Toyota Motor North America and conducted as a collaborative effort between its members and those from UCSB. We sincerely appreciate the members of the Human-AI Experience (HAX) Lab and other students for their participation in data collection experiments and their valuable feedback. We also extend our gratitude to the members of Toyota Motor North America, Toyota Motor Corporation, and Toyota Central Research and Development Labs Inc. for their insightful discussions and contributions. Additionally, the human subject study conducted in this research was reviewed and approved by the UCSB Human Subjects Committee under protocol number 14-22-0646.
@@ Appendix C Ablation study
Appendix D Additional experiments
D.1 Testing on GIMO
While the main paper utilized real-world trajectory data (LocoReal) as the test data, we conducted an additional experiment to further validate the contribution of LocoVR using an alternative test dataset. Given that THOR-MAGNI has fewer variational scenes, we selected GIMO as the test data. In this experiment, GIMO was divided into 70% for training, 15% for validation, and 15% for testing, ensuring that the scenes in the training, validation, and test sets were mutually exclusive. Additionally, to mitigate potential bias in the test data, we performed five random splits to generate five different datasets and averaged the results. Tables 8 through 10 illustrate the evaluation results for global path planning, trajectory prediction, and goal prediction, respectively. The evaluation results demonstrate that the model trained with LocoVR consistently outperforms those trained on other datasets. This superior performance is attributed to the enhanced generalization capabilities provided by LocoVR’s extensive coverage of scenes and trajectories.
@@ D.1 Testing on GIMO
D.2 Influence of scene information - Binary obstacle - Height - Semantics
To explore the potential utility of the semantic and height information included in the scene map, we conducted a small experiment to evaluate how replacing binary obstacle maps with 3D height maps and semantic maps affects performance. Table11 presents the results of the global path prediction task using the UNet+A* model. Each model was trained and tested on LocoVR with binary maps, height maps, and semantic maps, over three trials. As shown in the table, the models trained with height and semantic maps clearly outperformed those trained with binary maps. Although we do not yet have a detailed analysis of these findings, they potentially suggest that human trajectories could be influenced by object attributes inferred from height and semantic information. For instance, participants might unconsciously maintain a distance from movable objects, such as chairs or doors, or adjust their trajectories based on the visual clearance provided by different object types. For example, walls, kitchen counters, and low tables offer varying degrees of vision clearance, with lower clearance potentially exerting subtle psychological pressure on trajectory planning. A detailed analysis on influence of variational scene information on the human trajectories could provide valuable insights from the perspectives of cognitive and behavioral sciences.
Figure 6: Examples of binary (obstacle) map, height map, and semantic map
Table 11: Global path prediction with binary/semantic/height maps - ADE (Average Displacement Error) between predicted and ground-truth time-series trajectories. The table reports averaged performance in all of the trajectories over three trials $\pm$ SD
$Method$ $0m\leq d\leq 3m$ $3m\leq d\leq 6m$ $6m\leq d$
Binary map 0.138$\pm$0.0006 0.183$\pm$0.0024 0.286$\pm$0.0113
Semantic map 0.137$\pm$0.0004 1.70$\pm$0.0046 0.216$\pm$0.0278
Height map 0.136$\pm$0.0011 0.165$\pm$0.0068 0.201$\pm$0.0219
@@ F.4 Alignment of VR and the real space
F.5 Avatar calibration
We employed inverse kinematics software, FINAL-IK to reconstruct avatar motion from tracked body motions on the head/waist/hands/feet. Due to the performance of FINAL-IK, the avatar motions may occasionally display unnatural joint movements. However, these slight inaccuracies in body motion do not impact the contribution of our experiment, as our focus is on room-scale human dynamics rather than fine-grained body movements. Our dataset currently includes raw data from sparse motion trackers, which is highly accurate (within a few millimeters). For users requiring precise avatar motion, applying state-of-the-art IK algorithms to the raw tracker data would reconstruct more accurate avatar movements than those displayed in our video.
@@ Appendix H Data augmentation
For the global path prediction, the time window with a length of 10.0s (1.0s for past trajectory and 9.0s for future trajectory) slides with an interval of 0.67s to extract trajectory segments from the raw trajectory. The duration of 10.0s is chosen to encompass the majority of trajectories in the dataset (as shown in Figure 13), while a time step of 0.67 seconds is sufficient to capture human movement at speeds of up to 3.0 m/s. A part of the trajectory segment that is out of the raw trajectory is padded with the last value observed in the raw trajectory.
For the trajectory prediction, the time window with a length of 4.0s (1.0s for past trajectory and 3.0s for future trajectory) slides with an interval of 0.67s to extract trajectory segments from the raw trajectory. We assume that 3.0s is an appropriate time range for future trajectory prediction, as this task is non-goal-conditioned and affected by rapidly increasing positional uncertainty over time. Note that all the trajectory segments are within the raw trajectory because the task focuses on trajectory prediction in local areas on the way to the goal.
For the goal prediction, the time window with a length of 6.0s (for past trajectory) slides with an interval of 0.67s to extract trajectory segments from the raw trajectory. We set 6.0 seconds as the maximum length for past observations, as the primary objective is to predict the goal before it is reached, making long observation periods unnecessary for evaluating goal prediction performance. A part of the trajectory segment that is out of the raw trajectory is padded with the initial value observed in the raw trajectory.